OpenStudy (anonymous):

multiply and simplify 3b^4/11q^2 * 121q^8/9b

OpenStudy (ash2326):

We have \[ \frac{3b^4}{11q^2} \times \frac{121q^8}{9b}\] Let's cancel the like terms \[\frac{3b^{\cancel 4 3}}{11 \cancel{q^2} } \frac{121q^{\cancel 8 6}}{9 \cancel b}\] We get now \[\frac{3b^3 \times 121 q^6}{11 \times 9}\] Now \[\frac{\cancel 3b^3 \times \cancel {121} 11 q^6}{\cancel {11} \times \cancel 9 3}\] we get \[\frac{11b^3 q^6}{3}\]